OpenStudy (anonymous):

find dy/dx if 2y-xy^3=8

OpenStudy (ranga):

Do implicit differentiation first.

OpenStudy (anonymous):

is this first step correct? 2dy/dx-(dy/dx y^3+x3xy^2 dy/dx)=0

OpenStudy (tkhunny):

?? Looks like you're a little happy with the dy/dx. It's all about the Chain Rule! \(\dfrac{d}{dx}y = \dfrac{dy}{dx}\) - You got this piece. \(\dfrac{d}{dx}(-xy^{3}) = (-x)\dfrac{d}{dx}y^{3} + y^{3}\dfrac{d}{dx}(-x) = (-x)(3y^{2})\dfrac{dy}{dx} + y^{3}(-1)\)

OpenStudy (anonymous):

thank you!

OpenStudy (ranga):

After you do the implicit differentiation on the left equate it to 0 as the right hand side is a constant which when differentiated will yield zero. Then gather up all the dy/dx terms on the left and express it as dy/dx = some expression. I prefer to use the notation y' to dy/dx in implicit differentiation as it is less of a clutter.

OpenStudy (anonymous):

I understand the second part, but I'm confused on the implicit differentiation (I'm making corrections on a test I didn't very poorly on). Can you please show me how?

OpenStudy (anonymous):

*did very poorly on, sorry

OpenStudy (ranga):

2y - xy^3 = 8 The first term is 2y. Differentiate it with respect to y. It will be 2dy/dx or 2y' The next term is xy^3 (we will take care of the negative sign later on) Differentiate xy^3. Apply product rule. First term * derivative of second + second term * derivative of first. x * (3(y^2)y') + (y^3)(1) = 3x(y^2)y' + y^3 Put them together: 2y' - (3x(y^2)y' + y^3) = 0 Gather y' terms together. y'(2 - 3xy^2) - y^3 = 0 y'(2 - 3xy^2) = y^3 y' = y^3 / (2 - 3xy^2)

OpenStudy (ranga):

The second line should be: Differentiate it with respect to x (not y)

OpenStudy (ranga):

In the following examples we are differentiating with respect to x: Derivative of x is 1 Derivative of x^2 is 2x Derivative of x^7 is 7x^6 Derivative of y is dy/dx (or y') Derivative of y^2 is (2y)dy/dx = 2yy' Derivative of y^7 is 7(y^6)dy/dx = 7(y^6)y'

OpenStudy (ranga):

Note: Derivative of y^7 with respect to y is 7y^6 But Derivative of y^7 with respect to x is 7y^6 * (dy/dx) (chain rule)

OpenStudy (anonymous):

Thank you very much!! This was very helpful!

OpenStudy (ranga):

you are welcome. Glad to be able to assist.

OpenStudy (nincompoop):

\[x^n = nx^{n-1}\]